Welcome to
Azur Electronics


April 2017

The HP 3763A Error Detector on initial testing failed the Operator's Checks. The Adjustments including power rails are ok. Converted the front panel connections: Clock Input and Input A from 75Ω to 50Ω.  The Trigger Output is already 50Ω. Front panel connections: Clock Output; Count Gate Input; Sync Input are all ECL. Started working through the Troubleshooting procedures.

Local and Interval Times
Interval Time should start up as 00 00. Tricky fault turned out to be a failed buffer/driver output A14 IC5 holding the bus low.

Counter Gate Output on Rear Panel had no output. Fuse  A8 F2 replaced.

None of the odd numbered Word switches working. Flip-Flop A7 IC9 had failed.

Testing A14 PCB
May 2017

to be continued